A baseball is thrown straight upward with a speed of .
a. How long will it rise? [3]
b. How high will it rise? [3]
c. How long after it leaves the hand will it return to the starting point?[5]
1
Expert's answer
2020-06-01T14:26:16-0400

Solution.

"v;"

a. "g_y=\dfrac{v_y-v_{0y}}{t}\implies t=\dfrac{vy-v_{0y}}{g_y};"

"v_y=0, t=\dfrac{-v_{0y}}{g_y};"

b. "h_y=\dfrac{vy^2-v_{0y}^2}{2g_y};"

"v_y=0, h_y=\dfrac{-v_{0y}^2}{2g_y};"

c. "t_{max}=2\sdot(-\dfrac{v_{0y}}{g_y})=-\dfrac{2v_{0y}}{g_y};"

Answer: a. "t=\dfrac{-v_{0y}}{g_y};"

b. "h_y=\dfrac{-v_{0y}^2}{2g_y};"

c. "t_{max}=-\dfrac{2v_{0y}}{g_y};"